# GRAFT: Gradient-Aware Fast MaxVol Technique for Dynamic Data Sampling

### v1

@@ 6 Conclusion
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the paper’s contributions, including the proposed GRAFT framework for efficient data subset selection, its gradient-aware sampling mechanism, and its benefits in reducing computational cost and environmental impact. The claims are supported by theoretical analysis and experimental results across multiple datasets.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper includes a dedicated "Limitations and Future Work" section (Section 5.1) that discusses several limitations, such as reliance on the quality of low-rank feature extraction, computational overhead from dynamic rank adjustment, and sensitivity to batch size and feature extraction methods.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The paper provides theoretical results (Theorems 1-4, Lemmas 1-3) with clear assumptions and complete proofs in the main text and Appendix A. The proofs are mathematically rigorous and well referenced.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The paper provides extensive experimental details, including datasets used (CIFAR-10, CIFAR-100, FashionMNIST, TinyImageNet, IMDB, DermanMNIST), model architectures (ResNet-18, ResNeXt29, MobileNetV2, BERT), hyperparameters (learning rates, batch sizes, weight decay), and evaluation metrics. The Appendix includes implementation details.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: The paper states that "The code implementation for GRAFT will be provided after acceptance," but no code or data is currently available in the submitted version.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The paper provides comprehensive experimental settings, including optimizer details (SGD with Nesterov momentum), learning rate schedules (CosineAnnealing), batch sizes, and data fractions used for subset selection. All details are in the Appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: The paper reports mean and standard deviation over multiple runs (e.g., Tables 2-3, Figure 5) and includes statistical significance tests (p-values) where applicable.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The paper specifies hardware (Tesla V100-SXM2 GPU, Intel Xeon Gold CPU), memory (16GB), and execution times for key experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
[…]

### v2

@@ 6 Conclusion
Supplementary Material
7 Notations
Table 6: Summary of notation used throughout the paper
Symbol Description Dim./Type
$\mathcal{X}=\{x_{i}\}_{i=1}^{n}$ Full training dataset of $n$ samples set, $|\mathcal{X}|=n$
$x_{i}$ $i$-th training example $\mathbb{R}^{d}$
$n$ Number of training samples scalar
$d$ Input-feature dimension scalar
$X\in\mathbb{R}^{m\times n}$ Data matrix used in Fig. 1 matrix
$B$ Number of mini-batches per epoch scalar
$X_{i}$ $i$-th mini-batch subset of $\mathcal{X}$
$K$ Mini-batch size scalar
$T$ Total training iterations scalar
$S$ Subset-refresh interval scalar
$\Theta$ Trainable model parameters vector / tensor
$L(\Theta;x)$ Loss on a single sample scalar
$\nabla_{\Theta}L(\Theta;x)$ Per-sample gradient $\mathbb{R}^{|\Theta|}$
$\bar{\mathbf{g}}_{i}$ or $\bar{g}$ Mean gradient of batch $X_{i}$ $\mathbb{R}^{|\Theta|}$
$g_{r}$ Gradient of $r$-th selected sample $\mathbb{R}^{|\Theta|}$
$G\in\mathbb{R}^{d\times K}$ Gradient matrix of a full batch matrix
$\mathbf{R}=\{R_{1},\dots,R_{c}\}$ Candidate subset sizes (ranks) set
$R$ A chosen rank / subset size scalar
$R^{\star}$ Rank minimising projection error scalar
$\mathcal{S}_{i}^{R}$ Indices of $R$ samples selected from $X_{i}$ index set
$\mathcal{S}_{i}$ Final index set for batch $i$ after rank search index set
$\mathcal{S}^{t}$ Aggregate subset used at iteration $t$ index set
$f(\cdot)$ Feature-extraction mapping (e.g. SVD encoder) function
$\mathbf{V}=f(X_{i})\in\mathbb{R}^{K\times R}$ Low-rank feature matrix for batch $X_{i}$ matrix
$\text{fastmax-volume}(\mathbf{V},R)$ Fast MaxVol row-selection routine algorithm
$\mathbf{G}_{R}$ Gradient matrix restricted to $\mathcal{S}_{i}^{R}$ $\mathbb{R}^{d\times R}$
$d_{R}$ Projection error $\|\bar{g}-\mathbf{G}_{R}\mathbf{G}_{R}^{\dagger}\bar{g}\|_{2}^{2}$ scalar
$P_{R}$ Orthogonal projector onto $\operatorname{span}(\mathbf{G}_{R})$ matrix
$\epsilon$ Tolerance on gradient-projection error scalar
$\sigma_{R+1}$ (R+1)(R\!+\!1)-st singular value of batch matrix scalar
$p=[p_{1},\dots,p_{R}]$ Row indices returned by MaxVol vector
II Number of batches aggregated at refresh scalar
$\langle\hat{g},g_{R}^{t}\rangle$ Cosine similarity between epoch and subset gradients scalar
$E$ Estimated CO2 emissions kg CO2
$\Phi_{\text{acc}}(E)$ Accuracy fidelity coefficient vs. emissions function
$\Psi(f)$ Accuracy vs. trained data fraction $f$ function
$\lambda,E_{0},H$ Parameters of the exponential efficiency curve scalars
8 Extended Algorithm
Gradient-Aligned Sampling periodically (every $S$ iterations) scans each training mini-batch to find the smallest subset of examples whose gradients closely span the full-batch gradient it first averages per-example gradients to obtain $\bar{\mathbf{g}}_{i}$, then, for each candidate rank $R_{i}$, uses a fast max-volume routine on the batch’s feature matrix to pick $R_{i}$ exemplars, forms their gradient matrix, and computes the reconstruction error $d_{R_{i}}$; the rank $R^{\star}$ that minimizes this error yields the chosen indices $\mathcal{S}_{i}$, and aggregating over batches gives the active subset $\mathcal{S}^{t}$ for that period. In iterations between renewals, the method simply reuses the previous subset, and all model parameters are updated using only the currently active $\mathcal{S}^{t}$, thus retaining the stability of large-batch gradients while evaluating far fewer samples.
Output: Sampled data subset indices $\mathcal{S}^{t}$
begin
 for $t=1,\ldots,T$ do
    Stage 1: Subset Selection
    if $t\bmod S==0$ then
       for each batch $\mathcal{X}_{i}\subset\mathcal{X}$ do
          for $R_{i}\in\mathbf{R}$ do
             $\mathcal{S}_{i}^{r}=\text{fastmax-volume}(\mathbf{V}_{i},R_{i})$
              Gradient error $d_{R_{i}}=\|\bar{\mathbf{g}}_{i}-\mathbf{G}_{R_{i}}\mathbf{G}_{R_{i}}^{\dagger}\bar{\mathbf{g}}_{i}\|_{2}^{2}$
           Optimal rank $R^{*}=\arg\min_{R_{i}}\{d_{R_{i}}\}_{i=1}^{|\mathbf{R}|}$
           Store the selected indices $\mathcal{S}_{i}=\mathcal{S}_{i}^{R^{*}}$
        Aggregate all subsets $\mathcal{S}^{t}=\{\mathcal{S}_{i}\}_{i=1}^{I}$
    else
         Retain the previous subset $\mathcal{S}^{t}=\mathcal{S}^{t-1}$
    Stage 2: Model Update
     Update all model parameters using $\mathcal{S}^{t}$
Algorithm 2 Training with Gradient-Aligned Sampling
9 Detailed Proofs
Remark  1 (Gradient Approximation Guarantee) Let $A\in\mathbb{R}^{K\times M}$ be a batch matrix, and let $V=U_{R}\in\mathbb{R}^{K\times R}$ denote the matrix of the top-$R$ left singular vectors of $A$. Let $S\subset[K]$, $|S|=R$, be the subset of rows selected by the MaxVol algorithm applied to $V$. If the gradient map $g(x)=\nabla_{\theta}L(\theta;x)$ is $L_{g}$-Lipschitz continuous, then
$\left\|\nabla_{\theta}L(\theta;A)-\nabla_{\theta}L(\theta;A(S,\cdot))\right\|_{2}\leq\frac{K}{R}L_{g}\sigma_{R+1},$
where $\sigma_{R+1}$ is the $(R+1)$-th singular value of $A$.
Proof.
Consider the singular value decomposition (SVD) of $A$, given by $A=U\Sigma V^{\top}$. Denote by $A_{R}=U_{R}\Sigma_{R}V_{R}^{\top}$ the rank-$R$ truncation of $A$, and let the residual matrix be $E\mathrel{\mathop{\ordinarycolon}}=A-A_{R}$, so that $\|E\|_{2}=\sigma_{R+1}$. By construction, $A_{R}$ captures the dominant $R$-dimensional subspace of $A$.
[…]
Since $A_{R}$ shares the same left singular vectors as $V$, these properties allow us to express any row $i$ of $A_{R}$ as a linear combination of the sampled rows:
$A_{R}(i,\mathrel{\mathop{\ordinarycolon}})=T_{i,\mathrel{\mathop{\ordinarycolon}}}\,A_{R}(S,\mathrel{\mathop{\ordinarycolon}}),\qquad\forall\,i\in[K].$ (2)
Consider now the gradient map $g(x)=\nabla_{\theta}L(\theta;x)$, assumed to be $L_{g}$-Lipschitz. For each row $i$, define the interpolated gradient $)\hat{g}_{i}\mathrel{\mathop{\ordinarycolon}}=\sum_{j\in S}T_{ij}g(A(j,\mathrel{\mathop{\ordinarycolon}}$)). Using the reconstruction property (2) and the Lipschitz continuity, we obtain
$\|g(A(i,\mathrel{\mathop{\ordinarycolon}}))-\hat{g}_{i}\|_{2}\leq L_{g}(1+\|T_{i,\mathrel{\mathop{\ordinarycolon}}}\|_{1})\|E(i,\mathrel{\mathop{\ordinarycolon}})\|_{2}\leq L_{g}(1+R)\sigma_{R+1},$
where we have used the bounds $\|T_{i,\mathrel{\mathop{\ordinarycolon}}}\|_{1}\leq R$ and $\|E(i,\mathrel{\mathop{\ordinarycolon}})\|_{2}\leq\sigma_{R+1}$.
To bound the overall gradient approximation error, consider the average gradient over all rows, $)\bar{g}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{K}\sum_{i=1}^{K}g(A(i,\mathrel{\mathop{\ordinarycolon}}$)), and the average over the selected subset, $)g_{S}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{R}\sum_{j\in S}g(A(j,\mathrel{\mathop{\ordinarycolon}}$)). Expanding $\bar{g}-g_{S}$ in terms of the interpolated gradients and leveraging the properties of the interpolation matrix, we find that the terms involving $T$ sum to zero, leaving
$\|\bar{g}-g_{S}\|_{2}\leq\frac{1}{K}\sum_{i=1}^{K}L_{g}(1+R)\sigma_{R+1}\leq\frac{K}{R}L_{g}\sigma_{R+1},$
where the final bound uses the fact that $K\geq R$ and $1+R\leq 2R$ for practical choices of $R$.
Finally, recognizing that $\bar{g}=\nabla_{\theta}L(\theta;A)$ and $)g_{S}=\nabla_{\theta}L(\theta;A(S,\mathrel{\mathop{\ordinarycolon}}$)) completes the proof:
$\left\|\nabla_{\theta}L(\theta;A)-\nabla_{\theta}L(\theta;A(S,\cdot))\right\|_{2}\leq\frac{K}{R}L_{g}\sigma_{R+1}.$
Theorem  1 (Convergence via Gradient-Aligned Subspace Sampling) Let $\bar{g}\in\mathbb{R}^{d}$ denote the average gradient over a batch of $K$ samples, and let $G_{R}\in\mathbb{R}^{d\times R}$ be a collection of $R$ gradients selected via MaxVol, spanning a full-rank subspace $\mathcal{S}_{R}$. If the projection error $\|\bar{g}-\text{Proj}_{\mathcal{S}_{R}}(\bar{g})\|_{2}\leq\varepsilon$, then gradient descent using the projected direction converges to a stationary point under standard assumptions [16].
Assumptions. Suppose the loss function $L$ is $C$-smooth, i.e., for all $\Theta,\Theta^{\prime}$,
$\|\nabla L(\Theta)-\nabla L(\Theta^{\prime})\|_{2}\leq C\|\Theta-\Theta^{\prime}\|_{2}.$
Assume also that the gradients are bounded, $\|\nabla L(\Theta)\|_{2}\leq G$ for all $\Theta$, and the projection error satisfies $\|\bar{g}-g_{\mathrm{proj}}\|_{2}\leq\varepsilon$, where $g_{\mathrm{proj}}=\text{Proj}_{\mathcal{S}_{R}}(\bar{g})$.
Then, for the gradient descent updates $\Theta_{t+1}=\Theta_{t}-\eta g_{\mathrm{proj}}^{(t)}$, the iterates converge to a stationary point where $\|\nabla L(\Theta)\|_{2}\leq\varepsilon G$.
Proof.
By $C$-smoothness of $L$, the loss after a gradient step satisfies
$L(\Theta_{t+1})\leq L(\Theta_{t})-\eta\langle\nabla L(\Theta_{t}),g_{\mathrm{proj}}^{(t)}\rangle+\frac{C\eta^{2}}{2}\|g_{\mathrm{proj}}^{(t)}\|_{2}^{2}.$
Since the projection error is at most $\varepsilon$, we have
$\langle\nabla L(\Theta_{t}),g_{\mathrm{proj}}^{(t)}\rangle=\langle\nabla L(\Theta_{t}),\bar{g}^{(t)}\rangle-\langle\nabla L(\Theta_{t}),\bar{g}^{(t)}-g_{\mathrm{proj}}^{(t)}\rangle\geq\|\nabla L(\Theta_{t})\|_{2}^{2}-\varepsilon G,$
where we use Cauchy–Schwarz and the gradient bound $\|\nabla L(\Theta_{t})\|_{2}\leq G$.
Plugging this estimate into the smoothness inequality and recalling $\|g_{\mathrm{proj}}^{(t)}\|_{2}\leq G$, we obtain
$L(\Theta_{t+1})\leq L(\Theta_{t})-\eta\left(\|\nabla L(\Theta_{t})\|_{2}^{2}-\varepsilon G\right)+\frac{C\eta^{2}G^{2}}{2}.$
Summing this inequality over $t=0,\dots,T-1$ and dividing by $T$ gives
$\frac{1}{T}\sum_{t=0}^{T-1}\|\nabla L(\Theta_{t})\|_{2}^{2}\leq\frac{L(\Theta_{0})-L(\Theta^{*})}{\eta T}+\varepsilon G+\frac{C\eta G^{2}}{2},$
where $L(\Theta^{*})$ denotes the minimum loss. Setting $\eta=1/\sqrt{T}$ ensures the first and last terms vanish as $T\to\infty$, leaving the bound $\|\nabla L(\Theta_{t})\|_{2}\leq\varepsilon G$ in the limit. This establishes convergence to a stationary point determined by the projection error. ∎
Lemma 1 (Projection Error Bound) Let $\tilde{G}_{R}$ be an orthonormal basis for $G_{R}$. Then
Proof.
Since $\tilde{G}_{R}$ is orthonormal, the projection of $\bar{g}$ onto $\mathrm{span}(G_{R})$ is given by
$g_{\mathrm{proj}}=\tilde{G}_{R}\tilde{G}_{R}^{\top}\bar{g}.$
The residual $r=\bar{g}-g_{\mathrm{proj}}$ is orthogonal to $\mathrm{span}(G_{R})$. By the Pythagorean theorem,
$\|\bar{g}\|_{2}^{2}=\|g_{\mathrm{proj}}\|_{2}^{2}+\|r\|_{2}^{2}.$
Because $\tilde{G}_{R}$ is orthonormal, we have $\|g_{\mathrm{proj}}\|_{2}=\|\tilde{G}_{R}^{\top}\bar{g}\|_{2}$. Substituting this, we obtain
$\|r\|_{2}^{2}=\|\bar{g}\|_{2}^{2}-\|\tilde{G}_{R}^{\top}\bar{g}\|_{2}^{2}=\|\bar{g}\|_{2}^{2}\left(1-\left\|\frac{\tilde{G}_{R}^{\top}\bar{g}}{\|\bar{g}\|_{2}}\right\|_{2}^{2}\right),$
which completes the proof. ∎
Corollary  1 [Dynamic Rank Adjustment Ensures Convergence] If the rank $R$ is dynamically adjusted so that $\|\bar{g}-\tilde{G}_{R}\tilde{G}_{R}^{\top}\bar{g}\|_{2}^{2}\leq\varepsilon$ at every iteration, then GRAFT ensures convergence to a local minimum.
Proof.
The result follows directly from Theorem 1, since the gradient approximation error remains bounded by $\varepsilon$ at each step, independent of changes in $R$. Specifically, increasing $R$ refines the subspace $\mathrm{span}(G_{R})$ and reduces $\varepsilon$, while decreasing $R$ is permissible only if the new error still satisfies $\varepsilon\leq\varepsilon_{\mathrm{target}}$. In both scenarios, the condition $\|\bar{g}-g_{\mathrm{proj}}\|_{2}\leq\varepsilon$ is maintained, thereby ensuring convergence. ∎
9.1 Complexity Analysis of GRAFT
Theorem 2 (Projected-gradient convergence) Under (A1)–(A3), SGD with projected directions $G_{R}G_{R}^{\dagger}\,\bar{g}_{i}$ satisfies
$\min_{t\leq T}\ \mathbb{E}\bigl{\|}\nabla L(\Theta_{t})\bigr{\|}_{2}^{2}\;\leq\;O\!\left(\tfrac{1}{\sqrt{T}}\right)\;+\;O(\varepsilon^{2}).$
Proof.
By (A1), $L$ is $L$-smooth. With stepsizes $\{\eta_{t}\}$ and stochastic gradients $\bar{g}_{i}$ that are unbiased with bounded second moment (A2), the standard descent lemma gives
$\mathbb{E}\!\left[L(\Theta_{t+1})\right]\leq\mathbb{E}\!\left[L(\Theta_{t})\right]-\eta_{t}\,\mathbb{E}\!\left\langle\nabla L(\Theta_{t}),\,G_{R}G_{R}^{\dagger}\,\bar{g}_{i}\right\rangle+\frac{L\eta_{t}^{2}}{2}\,\mathbb{E}\!\left\|G_{R}G_{R}^{\dagger}\,\bar{g}_{i}\right\|_{2}^{2}.$
Decompose $\bar{g}_{i}=\nabla L(\Theta_{t})+\xi_{t}$ with $\mathbb{E}[\xi_{t}\mid\Theta_{t}]=0$. Write the projection error $e_{t}\mathrel{\mathop{\ordinarycolon}}=\bar{g}_{i}-G_{R}G_{R}^{\dagger}\bar{g}_{i}$ so that $\|e_{t}\|_{2}^{2}\leq\varepsilon^{2}$ by (A3). Then
$\left\langle\nabla L(\Theta_{t}),\,G_{R}G_{R}^{\dagger}\,\bar{g}_{i}\right\rangle=\left\langle\nabla L(\Theta_{t}),\,\bar{g}_{i}\right\rangle-\left\langle\nabla L(\Theta_{t}),\,e_{t}\right\rangle.$
Taking conditional expectation and using $\mathbb{E}[\xi_{t}\mid\Theta_{t}]=0$ yields
$\mathbb{E}\!\left[\left\langle\nabla L(\Theta_{t}),\,G_{R}G_{R}^{\dagger}\,\bar{g}_{i}\right\rangle\bigm{|}\Theta_{t}\right]\;\geq\;\|\nabla L(\Theta_{t})\|_{2}^{2}\;-\;\|\nabla L(\Theta_{t})\|_{2}\,\mathbb{E}\!\left[\|e_{t}\|\mid\Theta_{t}\right].$
By (A3), $\mathbb{E}[\|e_{t}\|\mid\Theta_{t}]\leq\varepsilon$, hence
$\mathbb{E}\!\left[\left\langle\nabla L(\Theta_{t}),\,G_{R}G_{R}^{\dagger}\,\bar{g}_{i}\right\rangle\right]\;\geq\;\mathbb{E}\!\left[\|\nabla L(\Theta_{t})\|_{2}^{2}\right]-\varepsilon\,\mathbb{E}\!\left[\|\nabla L(\Theta_{t})\|_{2}\right].$
For the quadratic term in (1),
$\mathbb{E}\!\left\|G_{R}G_{R}^{\dagger}\bar{g}_{i}\right\|_{2}^{2}\leq\mathbb{E}\!\left\|\bar{g}_{i}\right\|_{2}^{2}\leq\sigma^{2}+\mathbb{E}\!\left\|\nabla L(\Theta_{t})\right\|_{2}^{2},$
using nonexpansiveness of orthogonal projection and (A2). Substitute (2)–(3) into (1), rearrange, sum from $t=1$ to $T$, and use the standard stepsize choice $\eta_{t}=\eta/\sqrt{T}$ (or any schedule giving $\sum\eta_{t}=\Theta(\sqrt{T})$ and $\sum\eta_{t}^{2}=O(1)$). Telescoping the left side and bounding the sums on the right yields
$\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\!\left\|\nabla L(\Theta_{t})\right\|_{2}^{2}\;\leq\;O\!\left(\tfrac{1}{\sqrt{T}}\right)\;+\;O(\varepsilon^{2}),$
where the $O(1/\sqrt{T})$ term arises from the stochastic and smoothness terms and the $O(\varepsilon^{2})$ term collects the contribution induced by the projection bias in (2). Taking the minimum over $t\leq T$ gives the stated bound. ∎
Computational complexity
One iteration of the method runs in $O\!\left(KR^{2}\right)\;+\;O\!\left(|\text{Rset}|\,R\,d\right)$ time and uses $O(Kd+dR+R^{2})$ memory. The optional $O(Rd)$ term accounts for recomputing gradients on the selected $R$ samples when reported separately. The basis/feature refresh costs $O(KdR)+O((K{+}d)R^{2})$ only when executed and is amortized by its period. The per-iteration term is independent of $N$. Fast MaxVol scans the $K$ rows of the $K{\times}R$ feature matrix with rank-1 volume updates, yielding $O(KR^{2})$. The rank sweep evaluates a projection criterion (e.g., $\|\bar{g}-P_{r}\bar{g}\|_{2}$) for $r\in\text{Rset}$; forming $P_{r}\bar{g}$ is $O(rd)$ and sums to $O(|\text{Rset}|\,R\,d)$. If gradients are explicitly recomputed for the selected set, this adds $O(Rd)$. The refresh step (e.g., randomized range finding / blocked QR) is invoked periodically and amortized; all operations are mini-batch local, hence no factor depends on $N$. $\square$
Per-iteration and amortized costs are summarized below; symbols follow the notation already introduced.
Table 7: Per-iteration vs. amortized costs (batch size $K$, feature dim. $d$, active rank $R$, candidate ranks Rset). The refresh is performed periodically and amortized by its interval.
Operation Cost
Fast MaxVol on $K{\times}R$ $O(KR^{2})$
Projection/alignment sweep over $r\in\text{Rset}$ $O(|\text{Rset}|\,R\,d)$
(Optional) gradients for selected $R$ samples $O(Rd)$
(Periodic) basis/feature refresh $O(KdR)+O((K{+}d)R^{2})$
9.2 Additional Experimental Details
In this supplementary section, we provide a comprehensive set of additional experiments to further assess the generalizability and effectiveness of the proposed sampling approach across diverse machine learning tasks. First, we evaluate our method in the context of fine-tuning large language models (LLMs), demonstrating its scalability and impact in high-dimensional settings. We also include results on classical regression benchmarks to illustrate the method’s applicability beyond deep learning. Beyond empirical results, we investigate the convexity properties of the proposed sampling technique, analyzing how the sample selection mechanism influences the underlying loss landscape. This analysis provides theoretical justification for the observed stability and convergence behavior in our experiments. To offer a fair and transparent evaluation, we compare our approach against several widely used training and sampling strategies, including random sampling, core-set methods, and recent gradient-based subset selection algorithms. We report the standard performance metrics emissions by each method.
Collectively, these extended experiments and analyses aim to provide a deeper understanding of the strengths and potential limitations of the proposed sampling framework.
9.3 Experimental Details on Fine-Tuning BERT
In this experiment, we apply the proposed sampling method to the fine-tuning of large language models (LLMs), with a focus on practical NLP tasks. Specifically, we evaluate our approach on the IMDB sentiment analysis dataset by fine-tuning a distilled variant of BERT containing approximately 92 million parameters. Our sampling strategy, GRAFT, was executed on the embedding representations generated by the language transformer, enabling efficient selection of informative samples for gradient computation. The IMDB dataset is a well-established benchmark for text analytics, comprising 50,000 movie reviews labeled by sentiment. The dataset is split evenly, with 25,000 reviews allocated for training and 25,000 for testing. This setup allows us to thoroughly evaluate the generalization performance and efficiency of our sampling approach in a real-world NLP context.
Experimental Setting.
For fine-tuning, we employed a batch size of 100, a constant learning rate of $5\times 10^{-5}$, and a weight decay of 0.0001. The BERT model was fine-tuned for 30 epochs under two data regimes: the full dataset, and a GRAFT-selected subset generated using our GRAFT algorithm. For GRAFT, both standard and warm-start variants were tested, with data subset selection performed every 10 epochs. To assess the effect of sample efficiency, only 35% of the original training data was retained for each GRAFT selection cycle.
9.4 Extended Details and Experiments on Different Fractions of Data
In this section, we systematically compare the performance of various subset selection methods across different subset sizes (5%, 15%, 25%, and 35%). To ensure a fair comparison, we utilized the ResNeXt29_32x4d architecture [25] for all datasets, training each model for 200 epochs, except for TinyImageNet and Caltech256, where a ResNet-18 model was used due to dataset-specific considerations. All experiments were conducted with a fixed batch size of 200 and an initial learning rate of 0.1, using stochastic gradient descent (SGD) as the optimizer. For TinyImageNet and Caltech256, a smaller batch size of 100 was employed. A CosineAnnealing learning rate scheduler was consistently applied to adjust the learning rate during training. All models were trained from scratch on an NVIDIA Tesla V100-SXM2 GPU (16GB) with an Intel(R) Xeon(R) Gold CPU, except for TinyImageNet and Caltech256, which were trained on an NVIDIA A100-SXM4 GPU (40GB).
Power consumption and CO2 emissions were estimated using the methodology and tooling from Budennyy et al. [19]. Specifically, following the eco2AI framework, the instantaneous power usage of the computing hardware (GPU and CPU) is monitored throughout the training process. The framework records the real-time power draw (in watts) using hardware sensors or system queries, integrating these measurements over the training duration to obtain the total energy consumption (in kilowatt-hours, kWh) as
$E=\frac{1}{3600}\sum_{i=1}^{N}P_{i}\Delta t_{i},$ (2)
where $P_{i}$ is the instantaneous power at time step $i$, $\Delta t_{i}$ is the time interval since the last measurement, and $N$ is the total number of intervals. Note: In our main text, we report CO2 emissions using the standard formula $\mathcal{E}=P\times t\times I$, which assumes a constant average power $P$ over training duration $t$. The eco2AI library [19] also supports a more granular mode based on real-time power monitoring and numerical integration, yielding equivalent results in our single-GPU training regime. To convert total energy consumption into CO2 emissions, eco2AI multiplies $E$ by a region-specific carbon intensity factor $C$ (kg CO2 per kWh), typically reflecting the local energy mix. The resulting emissions are calculated as
[…]